My and d of this A vector has an x Need Help? My nature of this problem, do not use rounded intermediate values in your
Answer #1

14)

Magnitude = sqrt(27^2 + 36.4^2) = 45.32units

Direction, theta = arctan (36.4/-27) = - 53.43 degree

Direction = 180 - 53.43 = 126.57 degree counterclockwise from +ve x axis

Comment in case any doubt please rate my answer ....

According to the HOMEWORKLIB RULES I can answer one question at a time so please post other questions seperately ....